OpenStudy (anonymous):

simplify (11 + 4i)(-5 -9i)

OpenStudy (mathstudent55):

Use FOIL to start. Remember that \(i^2 = -1\).

OpenStudy (anonymous):

okay thanks.

OpenStudy (anonymous):

I got -55-99i-20-36i^2

OpenStudy (mathstudent55):

So far, you are almost correct. Just make sure the third term is -20i (add the i to what you have). Now use the fact that \(i^2 = -1\) to simplify the term \(-36i^2\).

OpenStudy (mathstudent55):

You should have this: -55 - 99i - 20i - 36i^2 Now you can do two things. 1. Combine -99i and -20i since they are like terms. 2. Change i^2 to -1, and simplify the last term.

OpenStudy (anonymous):

To check your answer, click on https://www.wolframalpha.com/input/?i=%2811+%2B+4i%29%28-5+-9i%29

OpenStudy (anonymous):

-55 - 119 ?

OpenStudy (anonymous):

-55 - 119 i

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

YW

OpenStudy (mathstudent55):

= -55 - 99i - 20i - 36i^2 The -99i term and the -20i term combine: = -55 - 119i - 36i^2 i^2 = -1, so we substitute: = -55 - 119i - 36(-1) = -55 - 119i + 36 = -19 - 119i That is the final answer.